# Opt2Skill: Imitating Dynamically-feasible Whole-Body Trajectories for Versatile Humanoid Loco-Manipulation

### v5

@@ I Introduction
Model-based methods that rely on an accurate model for control have been foundational in enabling humanoid skills. However, their reliability in multi-contact scenarios often limits them to behaviors with simple, pre-defined contacts. On the other hand, model-free RL that trains skills from scratch achieves robustness to complex contact modes even with environment uncertainties. However, it requires heavy reward tuning to regulate the policy behavior and to cross the sim-to-real gap. As a result, providing RL with a reference has become a mainstream approach. For example, learning loco-manipulation by imitating human motions has shown success on humanoid hardware [2, 3, 4, 5]. However, acquiring such data in high quality and for diverse tasks still poses a challenge. Furthermore, these datasets require motion retargeting, and the retargeted motions may not always be kinematically or dynamically feasible, limiting the robot’s ability to imitate the original motions effectively, and potentially losing behavior versatility to some extent. Other than using human data, complementing RL with model-based methods for whole-body control of legged robots is also a growing trend [6]. Along this line, learning dynamically-feasible whole-body motions from model-based methods to achieve diverse humanoid loco-manipulation tasks remains an open problem.
@@ II-A Model-Based Trajectory Optimization for Humanoids
Model predictive control (MPC) widely uses OCP for motion planning in an online receding horizon fashion. However, due to the computation constraint, MPC for humanoids predominantly uses various simplified models [10, 11], which result in conservative motions [12]. MPC incorporating full-body models [13, 14, 15, 16, 17] or kino-dynamic models [18, 19] still require high computational power and precise contact state feedback, which poses challenges for whole-body planning in real-time, contact-rich loco-manipulation tasks. Alternatively, TO can synthesize high-quality trajectories that are tracked by whole-body controllers. However, model-based whole-body controllers are sensitive to uncertainty in robot dynamics and environment contacts.
@@ II-B Reinforcement Learning for Humanoids
Reinforcement learning (RL) enables real-time whole-body control of humanoid robots, offering robustness to environmental uncertainty [20, 21, 22]. RL policies are trained through direct interactions with the environment, providing a straightforward training paradigm that learns from scratch without requiring pre-planned motions. However, such RL training often suffers from poor data efficiency. Furthermore, policies learned from scratch can produce unnatural motions [23] and require significant efforts in reward shaping.
@@
To address these challenges, RL-based motion imitation has been widely explored to enable humanoid loco-manipulation skills [2]. Compared with RL learned from scratch, motion imitation simplifies reward tuning and demonstrates improved data efficiency, yet results in natural motions. Therefore, there is a growing trend in complementing learning-based approaches with motion reference for whole-body control of legged robots  [6, 24].
@@
Motion references come from various sources. While such manually designed motion references can yield compelling behaviors [25, 21, 26], such motion references are often task-specific and do not generalize to diverse or versatile motions. Human motion capture data is another primary source for motion imitation [1], and has been successfully transferred to humanoid hardware via RL-based motion imitation [4, 27, 3], enabling both versatile teleoperation [5] and autonomous skills execution [28]. However, a significant embodiment gap remains between humans and humanoid robots, necessitating motion retargeting to adapt human movements to the robot’s morphology. So far, motion retargeting is largely restricted to upper-body manipulation tasks, and whole-body loco-manipulation retargeting often yields dynamically infeasible motions that require substantial data post-processing [5]. As such, the embodiment gap may limit the robot’s ability to faithfully imitate original whole-body motions and reduce behavioral versatility.
@@
To address these challenges, recent studies have begun using model-based TO to generate data for motion imitation. Model-based TO produces high-quality motion data by ensuring that the trajectory adheres to robot dynamics and other physical constraints, such as joint torque limits. TO-based motion imitation has been widely employed on quadrupeds [29, 6, 30]. For humanoid robots, various models have been leveraged for TO-based motion imitation, including single rigid-body model [31], centroidal dynamics model [32], kino-dynamic model [33], and full-order-dynamics model [34, 35]. However, most of these studies focus on only locomotion [36] rather than whole-body loco-manipulation. For loco-manipulation tasks, the high fidelity of the full-order model is beneficial as it eliminates the requirement for retargeting and is easier to obtain than compared to teleoperation. Furthermore, full-order motion data includes the joint torque, which is distinguished from other motion sources such as human motion capture or reduced-order motion. This work adopts a full-order-dynamics-based TO to guide RL, particularly for humanoid loco-manipulation tasks. We demonstrate that the joint torque information is critical for learning high-dimensional contact-rich skills. A closely related study is [30], which also uses torque information from TO for RL motion imitation. Our approach differs from [30] in several key aspects, including the complexity of the robot dynamics, the method used for tracking torque reference, and the extent of hardware validation.
@@ III Methods
Figure 2: Overall structure of the Opt2Skill framework. (a) We first generate structured, dynamically feasible reference trajectories using trajectory optimization with contact constraints, torque limits, and task-specific objectives. (b) Each trajectory contains joint angles, joint velocitis, body position, orientation, linear and angular velocities, as well as dynamic-relevant quantities such as joint torques and interaction forces. (c) These trajectories serve as supervision signals to train RL policies that predict joint-level targets, which are further tracked by a low-level PD controller. The resulting policies internalize control strategies grounded in model-based optimization while remaining reactive and robust to disturbances, sensor noise, and dynamics variability, enabling direct deployment to real hardware.
@@ III-A Whole-Body Trajectory Optimization
where $\mathbf{q}=\begin{bmatrix}\mathbf{q^{\rm b}};\mathbf{q^{\rm j}}\end{bmatrix}\in\mathbb{R}^{n_{q}}$ and $\mathbf{v}=\begin{bmatrix}\mathbf{v}^{\rm b};\dot{\mathbf{q}}^{\rm j}\end{bmatrix}\in\mathbb{R}^{n_{v}}$ are the generalized coordinates and velocities, partitioned into base ($\rm b$) and joint ($\rm j$) components. The base coordinates $\mathbf{q^{\rm b}}=\begin{bmatrix}\mathbf{p^{\rm b}};\bm{\theta}^{\rm b}\end{bmatrix}\in\mathbb{R}^{6}$ include base position and orientation, while the base velocities $\mathbf{v}^{\rm b}=\begin{bmatrix}\mathbf{\dot{p}^{\rm b}};\bm{\omega}^{\rm b}\end{bmatrix}\in\mathbb{R}^{6}\vspace{1mm}$ are linear and angular velocities in the world frame. $\mathbf{M(q)}$ is the joint-space mass matrix, and $\mathbf{C(q,v)}$ captures the nonlinear effects. $\mathbf{u}\in\mathbb{R}^{n_{j}}$ is the joint torque command, $\mathbf{J_{\rm c}(q)}$ is the contact Jacobian, and $\mathbf{F}_{\rm c}$ is the stacked contact reaction force vector. Given this whole-body dynamics, the TO formulation for generating a whole-body trajectory is written as:
$\underset{\mathbf{x},\mathbf{u}}{\min}$ $\sum_{k=0}^{N-1}\Biggl{(}\|\mathbf{y}[k]-\hat{\mathbf{y}}[k]\|^{2}_{Q}+\|\mathbf{u}[k]\|^{2}_{R}\Biggr{)}+$ (2a)
@@
[…]
@@
where the decision variables include the full-body state vector $\mathbf{x}=\begin{bmatrix}\mathbf{q};\dot{\mathbf{q}}\end{bmatrix}\in\mathbb{R}^{n_{q}+n_{v}}\vspace{1mm}$ and the joint torque $\mathbf{u}$. The task space variables $\mathbf{y}=\Phi(\mathbf{q})$ and $\hat{\mathbf{y}}$ are defined based on the specific task. In (2b), the hybrid dynamics constraint includes either the continuous whole-body dynamics (first row), or the impact dynamics (second row), where $\bm{\Lambda}$ denotes the contact impulse, and $\dot{\mathbf{q}}^{-}$ and $\dot{\mathbf{q}}^{+}$ are instantaneous velocities before and after the impact. Following (2b), the contact constraints (2c) is added for the stance or impact foot assuming a rigid contact, while a zero force is applied to the non-contact foot. Joint limit $\mathcal{J}$, torque limit $\mathcal{T}$, and friction cone $\mathcal{F}$ with coefficient $\mu$ are imposed as constraints in (2d)-(2e).
@@
To support robust policy learning across a wide range of loco-manipulation behaviors, we generate reference trajectories for diverse tasks, including walking, jumping, stair climbing, object pickup, and desk-object reaching. For each task, we systematically vary key motion parameters such as gait phase, contact mode, walking speed, foot clearance, center of mass position, and target object location. Leveraging the flexibility of our TO framework, we can efficiently sample these parameters to produce large datasets of motions. Unlike methods based solely on human demonstrations or heuristic design, our approach generates dynamically feasible trajectories, which contain critical dynamics-relevant properties such as joint torques and interaction forces that are often missing from kinematic demonstrations. Detailed task-specific definitions of $\hat{\mathbf{y}}$ and trajectory examples are presented in Sec. IV.
@@ III-B RL-Based Imitation of Dynamically Feasible Trajectories
III-B1 Problem Formulation
We formulate the control problem as an RL problem under a Markov Decision Process (MDP) framework, defined by the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma)$, with the state space $\mathcal{S}$, the action space $\mathcal{A}$, the transition dynamics $\mathcal{P}$, the reward function $\mathcal{R}$, and the discount factor $\gamma$. At each timestep $t$, the agent observes a state $s_{t}\in\mathcal{S}$, selects an action $a_{t}\sim\pi(a|s_{t})$, receives a reward $r_{t}$, and transitions to a new state $s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t})$.
In real-world deployments, the agent operates under partial observability due to sensor noise and unobservable environmental factors, resulting in a Partially Observable Markov Decision Process (POMDP). The policy $\pi(a|o_{t})$ must act based on noisy and incomplete observations $o_{t}\in\mathcal{O}$, which are incomplete projections of the true state $s_{t}$. Our goal is to learn an optimal policy $\pi$ that maximizes the expected discounted return $\mathbb{E}[\sum_{t=0}^{T}\gamma^{t}r_{t}]$, where the reward encourages accurate tracking of reference trajectories generated via TO.
We train the RL policy in the MuJoCo Playground simulator [38] using an Actor-Critic (AC) architecture optimized by Proximal Policy Optimization (PPO) [39]. Both the actor and critic networks are composed of a Multi-Layer Perceptron MLP with hidden sizes of $[512,512,256,256]$. All trainings are conducted on Nvidia RTX4090 GPUs.
III-B2 Imitation Policy
The observation space for the critic is defined as $\mathbf{o}_{\text{critic}}=\left[\mathbf{s}_{t}^{cp};\ \mathbf{s}_{t}^{cr}\right],$ where the privileged proprioception $\mathbf{s}_{t}^{cp}$ is $[\mathbf{p}^{b}_{t},\ \bm{\theta}^{b}_{t},\ \dot{\mathbf{p}}^{b}_{t},\ \bm{\omega}^{b}_{t},\ \mathbf{g}_{t},\ \mathbf{q}_{\text{hist}}^{\rm j},\ \dot{\mathbf{q}}^{\rm j}_{t},\ \mathbf{p}^{e}_{t},\ \mathbf{a}_{\text{hist}},\ \mathbf{{F}}^{\rm c}_{t},\ \mathbf{{u}}_{t},\ \mathbf{K}_{p},\ \mathbf{K}_{d}],$ which includes the humanoid body translation $\mathbf{p}^{b}_{t}$, orientation $\bm{\theta}^{b}_{t}$, linear velocity $\dot{\mathbf{p}}^{b}_{t}$, angular velocity $\bm{\omega}^{b}_{t}$, projected gravity $\mathbf{g}_{t}$ (a proxy for base orientation), a history of $N=10$ past motor joint positions sampled every $\delta=4$ timesteps, $\mathbf{q}_{\text{hist}}^{\rm j}=[\mathbf{q}_{t-1}^{\rm j},\mathbf{q}_{t-(1+\delta)}^{\rm j},\ldots,\mathbf{q}_{t-(1+(N-1)\delta)}^{\rm j}]$ (i.e., $50$ Hz sampling from a $200$ Hz control loop), motor joint velocities $\dot{\mathbf{q}}^{\rm j}_{t}$, end-effector position (relative to the torso) $\mathbf{p}^{e}_{t}$, a history of N past actions sampled with the same schedule, $\mathbf{a}_{\text{hist}}=[\mathbf{a}_{t-1},\mathbf{a}_{t-(1+\delta)},\ldots,\mathbf{a}_{t-(1+(N-1)\delta)}]$, contact force $\mathbf{{F}}^{\rm c}_{t}$, joint torque $\mathbf{{u}}_{t}$, and the PD gains $\mathbf{K}_{p}$, $\mathbf{K}_{d}$. The reference state $\mathbf{s}_{t}^{cr}$ is defined as $[\hat{\mathbf{p}}^{\rm b}_{t},\ \hat{\bm{\theta}}^{\rm b}_{t},\ \hat{\dot{\mathbf{p}}}^{\rm b}_{t},\ \hat{\bm{\omega}}^{\rm b}_{t},\ \hat{\mathbf{q}}_{t}^{\rm j},\ \hat{\dot{\mathbf{q}}}^{\rm j}_{t},\ \hat{\mathbf{p}}^{\rm e}_{t},\ \hat{\mathbf{{F}}}^{\rm c}_{t},\ \hat{\mathbf{{u}}}_{t}],$ where the hat $(\hat{\cdot})$ denotes reference trajectory information.
The observation space for the actor is defined as $\mathbf{o}_{\text{actor}}=\left[\mathbf{s}_{t}^{ap},\ \mathbf{s}_{t}^{ar}\right],$ where the noisy proprioception $\mathbf{s}_{t}^{ap}$ is $[\tilde{\dot{\mathbf{p}}}^{\rm b}_{t},\ \tilde{\bm{\omega}}^{\rm b}_{t},\ \tilde{\mathbf{g}}_{t},\ \tilde{\mathbf{q}}_{\text{hist}}^{\rm j},\ \tilde{\dot{\mathbf{q}}}^{\rm j}_{t},\ \mathbf{a}_{\text{hist}}],$ where the tilde $(\tilde{\cdot})$ indicates noisy sensor measurements. The actor also receives partial reference information $\mathbf{s}_{t}^{ar}$ = $[\hat{\dot{\mathbf{p}}}^{\rm b}_{t},\ \hat{\bm{\omega}}^{\rm b}_{t},\ \hat{\mathbf{q}}_{t}^{\rm j},\ \hat{\mathbf{{F}}}^{\rm c}_{t},\ \hat{\mathbf{{u}}}_{t}].$
The action space of the control policy represents offsets added to a default joint position (standing pose) to specify target positions for the $20$ actuated joints. These targets are fed into a PD joint torque controller, where the torque command is computed as $\mathbf{u}_{t}=\mathbf{K}_{p}(\mathbf{a}_{t}+\mathbf{{{q}}}^{\rm j}_{\text{dflt}}-\mathbf{q}^{\rm j}_{t})-\mathbf{K}_{d}\mathbf{{\dot{q}}}^{\rm j}_{t}$. Here, $\mathbf{{{q}}}^{\rm j}_{\text{dflt}}$ denotes the default standing joint positions, and $\mathbf{{{q}}}^{\rm j}_{t}$ and $\mathbf{{\dot{q}}}^{\rm j}_{t}$ are the measured joint positions and velocities, respectively. Our control policy runs at $200$ Hz, while the internal PD controller operates at $1$ kHz in simulation (during training) and $2$ kHz on hardware (during deployment).
@@ III-B2 Imitation Policy
Penalty Cost Action Rate $\|\mathbf{a}_{t}-2\mathbf{a}_{t-1}+\mathbf{a}_{t-2}\|_{2}^{2}$ $-0.05$
Torques $\|\mathbf{u}_{t}/\mathbf{u}_{\text{limit}}\|_{2}^{2}$ $-0.03$
Joint Acc. $\|\mathbf{{\ddot{q}}}^{\rm j}_{t}\|_{2}^{2}$ $-10^{-6}$
@@
Delays Action Delay Uniform $[0.0,0.02]\,\text{s}$

### v6

@@ I Introduction
Model-based methods that rely on an accurate model for control have been foundational in enabling humanoid skills. However, their reliability in multi-contact scenarios often limits them to behaviors with simple, pre-defined contacts. In constract, model-free RL that trains skills from scratch achieves robustness to complex contact modes even with environment uncertainties. However, it requires heavy reward tuning to regulate the policy behavior and to cross the sim-to-real gap. As a result, providing RL with a reference has become a mainstream approach. For example, learning loco-manipulation by imitating human motions has shown success on humanoid hardware [2, 3, 4, 5]. However, acquiring such data in high quality and for diverse tasks still poses a challenge. Furthermore, these datasets require motion retargeting, and the retargeted motions may not always be kinematically or dynamically feasible, limiting the robot’s ability to imitate the original motions effectively, and potentially losing behavior versatility to some extent. Other than using human data, complementing RL with model-based methods for whole-body control of legged robots is also a growing trend [6]. Along this line, learning dynamically-feasible whole-body motions from model-based methods to achieve diverse humanoid loco-manipulation tasks remains an open problem.
@@ II-A Model-Based Trajectory Optimization for Humanoids
Model predictive control (MPC) widely uses OCP for motion planning in an online receding horizon fashion. However, due to the computation constraint, MPC for humanoids predominantly uses various simplified models [10, 11, 12, 13], which result in conservative motions [14]. MPC that incorporates full-body models [15, 16, 17, 18, 19] or kino-dynamic models [20, 21] still require high computational power and precise contact state feedback, which poses challenges for whole-body planning in real-time, contact-rich loco-manipulation tasks. Alternatively, TO can synthesize high-quality trajectories that are tracked by whole-body controllers. However, model-based whole-body controllers are sensitive to uncertainty in robot dynamics and environment contacts.
@@ II-B Reinforcement Learning for Humanoids
Reinforcement learning (RL) enables real-time whole-body control of humanoid robots, offering robustness to environmental uncertainty [22, 23]. RL policies are trained through direct interactions with the environment, providing a straightforward training paradigm that learns from scratch without requiring pre-planned motions. However, such RL training often suffers from poor data efficiency. Furthermore, policies learned from scratch can produce unnatural motions [24] and require significant efforts in reward shaping.
@@
To address these challenges, RL-based motion imitation has been widely explored to enable humanoid loco-manipulation skills [2]. Compared with RL learned from scratch, motion imitation simplifies reward tuning and demonstrates improved data efficiency, yet results in natural motions. Therefore, there is a growing trend in complementing learning-based approaches with motion references for whole-body control of legged robots  [6, 25].
@@
Motion references come from various sources. While such manually designed references can yield compelling behaviors [26, 23, 27], they are often task-specific and do not generalize to diverse or versatile motions. Human motion capture data is another primary source [1], and has been successfully transferred to humanoid hardware via RL-based motion imitation [3, 28], enabling both versatile teleoperation [4] and autonomous skills [29]. However, a significant embodiment gap remains between humans and humanoids, necessitating nontrivial motion retargeting to adapt human movements to robot morphology. So far, motion retargeting is largely restricted to upper-body manipulation tasks, and whole-body loco-manipulation retargeting often yields dynamically infeasible motions that require substantial data post-processing [4]. As such, the embodiment gap may limit the robot’s ability to faithfully imitate original whole-body motions and reduce behavioral versatility.
@@
To address these challenges, recent studies have begun using model-based TO to generate data for motion imitation. Model-based TO produces high-quality motion data by ensuring that trajectories adhere to robot dynamics and other physical constraints, such as joint torque limits. TO-based motion imitation has been widely employed on quadrupeds [30, 6, 31]. For humanoids, various TO models have been leveraged, including single rigid-body model [32], centroidal dynamics model [33], kino-dynamic model [34], and full-order-dynamics model [35, 36]. However, most studies focus on only locomotion [37]. Instead, this work focuses particularly on humanoid loco-manipulation. To achieve this, our TO specifies not only the lower-body leg motions but also the upper-body contact sequence for object manipulation using full-order dynamics. Therefore, our reference trajectories are no longer restricted to periodic motions, highlighting the capability of learning versatile behaviors. Moreover, our work is the first to demonstrate hardware success in policies learned from full-order-dynamics-based TO on a humanoid robot. For loco-manipulation tasks, the high fidelity of the full-order model is beneficial as it eliminates retargeting and is easier to obtain compared to teleoperation. Furthermore, full-order motion data includes joint torques, distinguishing it from other sources such as human motion capture or reduced-order motion. We demonstrate that joint torque information is critical for learning high-dimensional contact-rich skills.
This work adopts a full-order-dynamics-based TO to guide RL. A closely related study is [31], which also uses TO-based torque supervision for RL. However, our approach differs in several important aspects: we focus on whole-body loco-manipulation using a $30$-DOF bipedal humanoid, while [31] focuses on quadrupedal locomotion with an $8$-DOF Solo88 robot; we generate trajectories using a full-order dynamics model, whereas they use a simplified rigid-body model; and our work includes extensive simulation and hardware evaluations across seven diverse tasks, along with quantitative analysis of sim-to-real transfer, the impact of torque/force reference information, and comparisons to other motion sources such as inverse kinematics and human demonstrations.
@@ III Methods
Figure 2: Overall structure of the Opt2Skill framework. (a) We first generate structured, dynamically feasible reference trajectories using trajectory optimization with contact constraints, torque limits, and task-specific objectives. (b) Each trajectory contains joint angles, joint velocities, body position, orientation, linear and angular velocities, and dynamic-relevant quantities such as joint torques and interaction forces. (c) These trajectories serve as supervision signals to train RL policies that predict joint-level targets, tracked by a low-level PD controller. The resulting policies internalize control strategies grounded in model-based optimization while remaining reactive and robust to disturbances, sensor noise, and dynamics variability, enabling direct deployment to real hardware.
@@ III-A Whole-Body Trajectory Optimization
where $\mathbf{q}=\begin{bmatrix}\mathbf{q^{\rm b}};\mathbf{q^{\rm j}}\end{bmatrix}\in\mathbb{R}^{n_{q}}$ and $\mathbf{v}=\begin{bmatrix}\mathbf{v}^{\rm b};\dot{\mathbf{q}}^{\rm j}\end{bmatrix}\in\mathbb{R}^{n_{v}}$ are the generalized coordinates and velocities, partitioned into base ($\rm b$) and joint ($\rm j$) components. The base coordinates $\mathbf{q^{\rm b}}=\begin{bmatrix}\mathbf{p^{\rm b}};\bm{\theta}^{\rm b}\end{bmatrix}\in SE(3)$ include base position and orientation represented as unit quaternion, while the base velocities $\mathbf{v}^{\rm b}=\begin{bmatrix}\mathbf{\dot{p}^{\rm b}};\bm{\omega}^{\rm b}\end{bmatrix}\in se(3)\vskip 2.84526pt$ are linear and angular velocities in the body frame. $\mathbf{M(q)}$ is the joint-space mass matrix, and $\mathbf{C(q,v)}$ captures the nonlinear effects. $\mathbf{u}\in\mathbb{R}^{n_{j}}$ is the joint torque command, $\mathbf{J_{\rm c}(q)}$ is the contact Jacobian, and $\mathbf{F}_{\rm c}$ is the stacked contact reaction force vector.
Given this whole-body dynamics, the TO formulation for generating a whole-body trajectory is written as:
$\underset{\mathbf{x},\mathbf{u}}{\min}$ $\sum_{k=0}^{N-1}\Bigl(\|\mathbf{y}[k]-\hat{\mathbf{y}}[k]\|^{2}_{Q}+\|\mathbf{u}[k]\|^{2}_{R}\Big)+$ (2a)
@@
$(\textit{\rm Dynamics})$ $\begin{cases}\mathbf{M}(\mathbf{q})\mathbf{\ddot{q}}+\mathbf{C}(\mathbf{q},\mathbf{\dot{q}})=\mathbf{B}\mathbf{u}+\mathbf{J_{\rm c}^{\top}}\mathbf{F_{\rm c}}\\ \mathbf{M}(\mathbf{q})\mathbf{\dot{q}}^{+}-\mathbf{M}(\mathbf{q})\mathbf{\dot{q}}^{-}=\mathbf{J_{\rm c}^{\top}}\bm{\Lambda}\end{cases}$ (2b)
$(\textit{\rm Contact})$ $\begin{cases}\mathbf{J}_{\rm c}\mathbf{\ddot{q}}+\mathbf{\dot{J}_{\rm c}}\mathbf{\dot{q}}=0\\ {\color[rgb]{0,0,0}\mathbf{J}_{\rm c}\mathbf{\dot{q}}^{+}=0}\end{cases}$ (2c)
$(\textit{\rm Limits})$ $\mathbf{q}^{\rm j}\in\mathcal{J},\mathbf{u}\in\mathcal{T}$ (2d)
@@
where the decision variables include the full-body state vector $\mathbf{x}=\begin{bmatrix}\mathbf{q};\dot{\mathbf{q}}\end{bmatrix}\in\mathbb{R}^{n_{q}+n_{v}}\vskip 2.84526pt$ and the joint torque $\mathbf{u}$. The task space variables $\mathbf{y}=\Phi(\mathbf{q})$ and $\hat{\mathbf{y}}$ are defined based on the specific task. In (2b), the hybrid dynamics constraint includes either the continuous whole-body dynamics (first row), or the impact dynamics (second row), where $\bm{\Lambda}$ denotes the contact impulse, and $\dot{\mathbf{q}}^{-}$ and $\dot{\mathbf{q}}^{+}$ are instantaneous velocities before and after the impact. Following (2b), the contact constraints (2c) is added for the stance or impact foot assuming a rigid contact, while a zero force is applied to the non-contact foot. Joint limit $\mathcal{J}$, torque limit $\mathcal{T}$, and friction cone $\mathcal{F}$ with coefficient $\mu$ are imposed as constraints in (2d)-(2e).
@@
To support robust policy learning across a wide range of loco-manipulation behaviors, we generate reference trajectories for diverse tasks, including walking, stair climbing, object pickup, drawing, and desk-object reaching. For each task, we systematically vary key motion parameters such as gait phase, contact mode, walking speed, foot clearance, center of mass position, and target object location. Leveraging the flexibility of our TO framework, we can efficiently sample these parameters to produce large datasets of motions. Unlike methods based solely on human demonstrations or heuristic design, our approach generates dynamically feasible trajectories, which contain critical dynamics-relevant properties such as joint torques and interaction forces that are often missing from kinematic demonstrations. Detailed task-specific definitions of $\hat{\mathbf{y}}$ and trajectory examples are presented in Sec. IV.
@@ III-B RL-Based Imitation of Dynamically Feasible Trajectories
During training, one trajectory is randomly sampled at the start of each episode to initialize the environment and serve as the reference for computing tracking rewards. We train separate policies for each loco-manipulation task, each using its own TO-generated trajectory dataset. At deployment, the policy is provided with partial references from a new offline-generated TO trajectory tailored to the specific test scenario, without reusing training data or running TO online.
While DDP (used within TO) produces dynamically feasible nominal trajectories, it is rarely employed directly for control due to modeling inaccuracies, contact uncertainties, and high computational cost in real-time settings. To address this, we leverage RL policies that track these trajectories and adapt to real-world disturbances during execution.
@@
Penalty Cost Action Rate $\|\mathbf{a}_{t}-2\mathbf{a}_{t-1}+\mathbf{a}_{t-2}\|_{2}^{2}$ 0.05-0.05
Torques $\|\mathbf{u}_{t}/\mathbf{u}_{\text{limit}}\|_{2}^{2}$ 0.03-0.03
Joint Acc. $\|\mathbf{{\ddot{q}}}^{\rm j}_{t}\|_{2}^{2}$ $10^{-6}$
@@
Delays Action Delay Uniform $0.0,0.02]\,\text{s}$
@@
III-B1 Problem Formulation
We formulate the control problem as an RL problem under a Markov Decision Process (MDP), defined by the tuple $\mathcal{M}=(\mathcal{S},\mathcal{A},\mathcal{P},\mathcal{R},\gamma)$, with state space $\mathcal{S}$, action space $\mathcal{A}$, transition dynamics $\mathcal{P}$, reward function $\mathcal{R}$, and discount factor $\gamma$. At each timestep $t$, the agent observes a state $s_{t}\in\mathcal{S}$, selects an action $a_{t}\sim\pi(a|s_{t})$, receives a reward $r_{t}$, and transitions to a new state $s_{t+1}\sim\mathcal{P}(\cdot|s_{t},a_{t})$. In real-world deployments, the agent operates under partial observability due to sensor noise and unobservable environmental factors, resulting in a Partially Observable Markov Decision Process (POMDP). The policy $\pi(a|o_{t})$ must act based on noisy and incomplete observations $o_{t}\in\mathcal{O}$, which are incomplete projections of the true state $s_{t}$. Our goal is to learn an optimal policy $\pi$ that maximizes the expected discounted return $\mathbb{E}[\sum_{t=0}^{T}\gamma^{t}r_{t}]$, where the reward encourages accurate tracking of reference trajectories generated via TO.
We train the RL policy in the MuJoCo Playground simulator [39] using an Actor-Critic (AC) architecture optimized by Proximal Policy Optimization (PPO) [40]. Both the actor and critic networks are composed of a Multi-Layer Perceptron MLP with hidden sizes of $[512,512,256,256]$. Each policy is trained for fewer than $300$ million steps, requiring under $7$ hours on a single NVIDIA RTX40904090 GPU.
III-B2 Imitation Policy
The observation space for the critic is defined as $\mathbf{o}_{\text{critic}}=\left[\mathbf{s}_{t}^{cp};\ \mathbf{s}_{t}^{cr}\right],$ where the privileged proprioception $\mathbf{s}_{t}^{cp}$ is $\mathbf{p}^{b}_{t},\ \bm{\theta}^{b}_{t},\ \dot{\mathbf{p}}^{b}_{t},\ \bm{\omega}^{b}_{t},\ \mathbf{g}_{t},\ \mathbf{q}_{\text{hist}}^{\rm j},\ \dot{\mathbf{q}}^{\rm j}_{t},\ \mathbf{p}^{e}_{t},\ \mathbf{a}_{\text{hist}},\ \mathbf{{F}}^{\rm c}_{t},\ \mathbf{{u}}_{t},\ \mathbf{K}_{p},\ \mathbf{K}_{d}],$ which includes the humanoid body translation $\mathbf{p}^{b}_{t}$, orientation $\bm{\theta}^{b}_{t}$, linear velocity $\dot{\mathbf{p}}^{b}_{t}$, angular velocity $\bm{\omega}^{b}_{t}$, projected gravity $\mathbf{g}_{t}$ (a proxy for base orientation), a history of $N=10$ past motor joint positions sampled every $\delta=4$ timesteps, $\mathbf{q}_{\text{hist}}^{\rm j}=[\mathbf{q}_{t}^{\rm j},\mathbf{q}_{t-\delta}^{\rm j},\ldots,\mathbf{q}_{t-(N-1)\delta}^{\rm j}]$ (i.e., $50$ Hz sampling from a $200$ Hz control loop), motor joint velocities $\dot{\mathbf{q}}^{\rm j}_{t}$, end-effector position (relative to the torso) $\mathbf{p}^{e}_{t}$, a history of N past actions sampled with the same schedule, $\mathbf{a}_{\text{hist}}=[\mathbf{a}_{t-1},\mathbf{a}_{t-(1+\delta)},\ldots,\mathbf{a}_{t-(1+(N-1)\delta)}]$, contact force $\mathbf{{F}}^{\rm c}_{t}$, joint torque $\mathbf{{u}}_{t}$, and the PD gains $\mathbf{K}_{p}$, $\mathbf{K}_{d}$. The reference state $\mathbf{s}_{t}^{cr}$ is defined as $\hat{\mathbf{p}}^{\rm b}_{t},\ \hat{\bm{\theta}}^{\rm b}_{t},\ \hat{\dot{\mathbf{p}}}^{\rm b}_{t},\ \hat{\bm{\omega}}^{\rm b}_{t},\ \hat{\mathbf{q}}_{t}^{\rm j},\ \hat{\dot{\mathbf{q}}}^{\rm j}_{t},\ \hat{\mathbf{p}}^{\rm e}_{t},\ \hat{\mathbf{{F}}}^{\rm c}_{t},\ \hat{\mathbf{{u}}}_{t}],$ where the hat (^)($\hat{\cdot}$) denotes reference trajectory information.
The observation space for the actor is defined as $\mathbf{o}_{\text{actor}}=\left[\mathbf{s}_{t}^{ap},\ \mathbf{s}_{t}^{ar}\right],$ where the noisy proprioception $\mathbf{s}_{t}^{ap}$ is $\tilde{\dot{\mathbf{p}}}^{\rm b}_{t},\ \tilde{\bm{\omega}}^{\rm b}_{t},\ \tilde{\mathbf{g}}_{t},\ \tilde{\mathbf{q}}_{\text{hist}}^{\rm j},\ \tilde{\dot{\mathbf{q}}}^{\rm j}_{t},\ \mathbf{a}_{\text{hist}}],$ where the tilde (~)($\tilde{\cdot}$) indicates noisy sensor measurements. The actor also receives partial reference information $\mathbf{s}_{t}^{ar}$ = $\hat{\dot{\mathbf{p}}}^{\rm b}_{t},\ \hat{\bm{\omega}}^{\rm b}_{t},\ \hat{\mathbf{q}}_{t}^{\rm j},\ \hat{\mathbf{{F}}}^{\rm c}_{t},\ \hat{\mathbf{{u}}}_{t}].$ The actor receives partial references to ensure generalization and avoid reliance on noisy or redundant inputs such as global base states or end-effector poses, while the critic has access to full privileged information for stable value estimation.
The action space of the control policy represents offsets from a default standing pose to specify target positions for the $20$ actuated joints. These targets are fed into a PD joint torque controller, which computes torque as $\mathbf{u}_{t}=\mathbf{K}_{p}(\mathbf{a}_{t}+\mathbf{{{q}}}^{\rm j}_{\text{dflt}}-\mathbf{q}^{\rm j}_{t})-\mathbf{K}_{d}\mathbf{{\dot{q}}}^{\rm j}_{t}$. Here, $\mathbf{{{q}}}^{\rm j}_{\text{dflt}}$ denotes the default standing joint positions, and $\mathbf{{{q}}}^{\rm j}_{t}$ and $\mathbf{{\dot{q}}}^{\rm j}_{t}$ are the measured joint positions and velocities. Our control policy runs at $200$ Hz, while the internal PD controller operates at $1$ kHz in simulation (during training) and $2$ kHz on hardware (during deployment). No explicit filtering is applied between the policy output and the PD control loop. In hardware, we use the same PD gains as in simulation and send torque commands directly to the robot.